how many liters of alkali must be added to 10 liters of water o be able to produce a solution that is 75% alkali
Answer by TimothyLamb(4379)

x = liters of alkali
---
x = 0.75(x + 10)
x = 0.75x + 7.5
x - 0.75x = 7.5
0.25x = 7.5
x = 7.5/0.25
---
answer:
x = liters of alkali = 30
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php